OpenStudy (anonymous):

The sum of one half of a number and 6 equals one third of the number, what is the number? a. -36 c. -24 b. -42 d. -48

hartnn (hartnn):

let the number be x. then, x/2 +6 = x/3 can u solve for x now??

OpenStudy (anonymous):

should the equation be 1\2x+6 =1/3x

hartnn (hartnn):

if u meant (1/2)x+6=(1/3)x then we both are correct .....

OpenStudy (anonymous):

i tried using the gcm and multiplying buy the reciprocals but i didnt get the correct answer

hartnn (hartnn):

multiply that whole equation by 6 first and tell me what u get.

OpenStudy (anonymous):

6/12 +36=6/18 ???

OpenStudy (pradipgr817):

a) -36

hartnn (hartnn):

but where did x go ?? (6/2)x+36=(6/3)x 3x+36=2x x=-36 that is correct,-36

OpenStudy (pradipgr817):

(x/2)+6=x/3 (x/2)-(x/3)=-6 x/6=-6 x=-36

OpenStudy (anonymous):

oooooo i see thnkz

hartnn (hartnn):

welcome :)